OpenStudy (anonymous):

I will give you a medal and be your fan What is 18 divided by 2/3? 1/12 3 12 27

OpenStudy (nikita2):

18/(2/3) = 27

OpenStudy (jack1):

18 / (2/3) = 18 times 3/2 = (18 * 3)/2 = 54/2 =27

OpenStudy (anonymous):

Wow,that's a great way to explain it

OpenStudy (zzr0ck3r):

I think at this point we are just giving answers. I have told you how to do this a few times now and it cant get any easier. try the next one by yourself.

OpenStudy (jack1):

;D, cheers